# Identification of phase correlations in Financial Stock Market Turbulence

### v1

@@ A Introduction
The Financial market is extremely reactive to changes in the market environment. A small information on underlying factors will lead to changes in market behavior and the pricing patterns. The information which are directly or indirectly related to the market affects the price movement and is vital for price discovery. Reflection of companies’ performance and overall dynamics of the market on price provides a foundation for the stakeholders to assess trends and try to discover the price to make informed decisions. The traces of price discovery go back to the 1900s when Bachelier talked about the application of probability theory to analyze the price and gain from options through his work “Theory of speculation”. Followed by Fama (1965) Fama (1970) laid the foundation for the efficient market hypothesis, arguing that prices are random and follow a random walk. He also identified three forms of market efficiency where past prices, public information, and private information play a role in determining the prices. But it is not possible to access all the available information, and no one is capable of accessing all the information, resulting in a random nature in price fluctuations and patterns. The study by Fama et al. (1969) revealed that the market prepares for splits, adjusts the prices accordingly, and is capable of fully incorporating the information, thus supporting the Efficient Market Hypothesis. Laffont and Maskin (1990) investigates whether the Efficient Market Hypothesis (EMH) holds under two conditions: first, where prices do not reveal private information, and second, where prices do. They analyze how large traders mask the information using Perfect Bayesian Equilibrium and comparative statistics. They found that the market reveals the information, but at the same time, the traders do not prefer to have the market as efficient. This indicates that the EHM is not applicable in imperfect competition when a large group of traders hides information, preventing the price from reflecting the private information that they hold. Mitchell and Mulherin (1994) also found that the news reported in Dow Jones & Company newspaper is positively correlated to the trading volumes and market returns of US securities, indicating a direct relation to the trading activity. Chan et al. (2008) studied information asymmetry between foreign and domestic shares in the Chinese stock market. Their analyses suggested the presence of information asymmetry between foreign and domestic shares in the Chinese stock market. The study by Urquhart and McGroarty (2016) found the traces of an efficient market on S&P 500 and EUROSTOXX 50 by applying AR-GARCH and Variance. Due to the impact of information on price discovery and fluctuations., Fama’s theory of Efficient Market Hypothesis is still relevant and is important to analyze the stages of market patterns and fluctuations to discover the quantitative aspect of the information in terms of price dynamics.
@@
The prices are impacted by all those direct or indirect factors, signaling that price fluctuations emerge from the changes and information in the market. Fama urged that the market incorporates all information in an efficient manner and the market are always efficient. Fama (1970). He also stated that the prices are always random and are capable of reflecting all available information, and the pricing patterns are always unpredictable. Due to the importance of the role of information on price discovery and pricing patterns, it makes investors difficult task to predict the market with certainty. But the development of different theories like Arbitrage Pricing Theory(APT) and Capital Assets Pricing Models (CAPM) helps to understand how various factors influence the asset pricing, providing an advantage to the arbitrage Ross et al. (1973)Fama and French (2004). Considering the information and the underlying factors, investors try to assess such trends and movements of price to make informed decisions, as such price reflects the overall performance of the listed companies and overall market dynamics. The rising stock prices indicate enhanced company performance, while declining prices indicate downward trends. Stock prices display cyclical movements that are influenced by the dynamic nature of micro as well as macroeconomic factors. followed by the reactions of investors to any economic factors can lead to variations in share prices, Martynova and Renneboog (2008). Each stock price exhibits unique cyclical patterns, which are studied separately or with the market. It is also important to observe that the cyclical patterns and behaviors of prices may have emerged from trends and autocorrelation, Aydogan and Booth (1988). To understand the market dynamics, one needs to analyze the pricing patterns and their cyclical fluctuations, which ultimately help to predict future price movements. Examining the cyclical components of time series data is important for understanding the market trends and forecasting the prices, supporting the interpretation of the movements, whether short or long period, Idrees et al. (2019). Forecasting the market depends on identifying the patterns in price sequences by deeply analyzing the historical prices to improve predictability, Cowles (1944).
@@
while discussing the price discovery. It is also equally important to recognize and understand that the patterns are not always uniform and are always influenced by the uncertainties and asymmetric information. Such interaction of the elements has offered the interest of examining the price assessment and prediction, ultimately leading an informed decision-making. A study by Malkiel (1989) suggested that to have more research on the dynamics of pricing to discover the true nature of pricing patterns, whether they are truly because of the nature of the market or a result of data exploitation. The fluctuations in the stock market are not always consistent and are impacted by various related factors, as well as events and information asymmetry, which causes turmoil in the market, signaling an unpredictable nature of the price movements. These regular ups and downs of the stock prices and continuous fluctuations of the movement of price are known as turbulence Wang et al. (2020) Mantegna and Stanley (1999). Such situations in the market create volatility, on which the cyclical patterns and movements of prices become difficult. The turbulence in the financial market is generally caused by the situation of information asymmetry and the high-energy events, and it is important to analyze the market turbulence and related factors. Engel et al. (2018) studied the major financial crisis, which caused the turbulence in the stock market and framed an early Warning signal using Markov switching models on such situations for Asian and European stock markets. Another study by Khoojine and Han (2020) developed the SPNAR stock prices network auto-regressive model to analyze the Chinese stock market during pre- and post-turbulence. Understanding such situations of uncertainty is essential for market analysis and helps to prevent such situations. Due to the flow of asymmetric information and the dynamic nature of the market, the market becomes volatile with rapid fluctuations, which makes it difficult to analyze and predict the market through the price pattern. So the examination of cyclical pattern ( rapid ups and downs) of the market is valuable to understand the nature of the market as turbulent or not. In this paper, we will analyze the Indian stock market, especially the National Stock Exchange of India, considering the factor of efficient market and the price dynamics through which the nature and state of turbulence within the NSE can be understood. The concepts and techniques of computational physics, especially the spectrum analysis, will be employed to understand the above-mentioned scenario, contributing a different approach for analyzing the market dynamics.
@@ A.1 Stock Market and Turbulence
The stock market is impacted by unpredictable events and the turbulence caused by such events needs to be understood to predict and understand the impact of such events on the stock market. The stock market is unpredictable and due to the rapid changes in environment and the events which affects the financial market leads to disturbances and situation of volatility and such unpredictable ups and downs of the stock market is termed to be known as the state of Turbulence. Turbulence in the stock market refers to rapid and significant increase and decrease in the prices, followed by the situation of unpredictability in the financial market. Turbulence signifies the span of extreme volatility and instability in the financial market which are characterized by the fluctuations and huge level of uncertainty Wang et al. (2020). The word turbulent is commonly used as the fluctuations in price and patterns which replicate the velocity fluctuations in the turbulence Mantegna and Stanley (1999).
@@
Turbulence can be viewed as a combination of various independent cycles/frequencies present in a system. Typically, in fluid turbulence, a fluid medium can sustain several independent frequencies at the same time. This happens as fluids cannot sustain strong shear forces and fluid elements try to thermalise through conduction as well as convection mechanism. When energy is pumped into a fluid medium, typically the energy resides in large wavenumbers or in long-time cycles of the fluid’s motion. However, this energy cascades into shorter wavenumbers or in short-time cycles soon. This cascading of energy into shorter scales is called the inertial range of turbulence. Finally, the energy is drained out in the form of heat, through tiny viscous scales of the system. Thus, typically a fully-developed turbulent fluid contains three clear scales – the first being the scale at which the energy is being injected into the system, the second is the inertial scale where the energy cascades to the shorter scales and the third is the viscous scale where the energy gets drained out from the medium in the form of heat/radiation. The second phase, that is, the inertial scale, plays the key role in developing the turbulence in a fluid. This inertial scale transfers the energy (/amplitude in the stock market) from large-scale vertices (/long-time cycles) to smaller vertices (/short-time cycles or fluctuations) through some nonlinear interactions. It has been found that in a fully developed turbulence, these nonlinear interactions between different modes (/cycles) actually retain information from both the amplitude as well as the phase of each of the modes (/cycles). Thus, merely an analysis using the energy spectrum falls short in understanding the actual nature of the turbulence present in the system. For example, inside a fully developed turbulence, we expect there will be no correlation between the individual phases, and thus each cycle will have its own unique phase information. In addition, if there is a huge correlation between the phases of individual cycles, one can conclude that the medium has not reached a fully developed turbulent phase. In this report, we describe a method to dig out this phase information from stock market data (using an integral transform method) and check for the quality or degree of turbulence in these individual data. This method is already in use to identify and quantify the level of turbulence in fluids as well as in plasmas Kumar et al. (2019a) Kumar et al. (2016). However, to the best of our knowledge, this is the first time we employ this technique to categorize turbulence in stock market data. Nussbaumer and Nussbaumer (1982)Kumar et al. (2019b)
@@ A.2 Market dynamics and Computational Approach
Due to the effectiveness of fundamental analysis to examine the stock market for long-term investors, it lacks effectiveness on short-term predictions as it considers a range of variables, including macroeconomic factors, industry dynamics, and company-specific elements, to determine the value of financial assets. However, this analysis can be complex, requiring consideration of multiple variables and significant time investment, making it more suitable for long-term investors. Graham et al. (1951), Graham and McGowan (2005). In contrast, due to the desire for quicker gains and faster analysis of market trends, technical analysis emerged. This approach posits that all available information is already reflected in stock prices, allowing analysts to focus on price trends using historical data to capitalize on market movements Hamilton (2006), Bishop Jr (1961), Schannep (2008), Benjamin (1942), Griffioen (2003). Technical analysis has experienced significant growth, particularly among short-term traders. Due to the complexity, noisy, chaotic, volatile, and vast amount of data that needs to be analyzed, machine-learning techniques have emerged and delivered superior results. The computational approach, generally a technical analysis, provides more meticulous outcomes and the ability to analyze big data. The data analysts nowadays are using computational and machine learning to analyze financial data. The stock market data are very complicated and non-stationary, making the traditional approaches typically inaccurate, and they treat stock data as a linear time series. The study by Kumar et al. (2021) considered such issues and reviewed the recent developments on the application of computational methods in the financial markets. Their study found that methods like Artificial Neural Networks, Support Vector Machines are intensively used in place of the traditional approach of data analysis. Similarly, Dunis et al. (2014) also provided different computational techniques that are used for analyzing the stock data. Sonkavde et al. (2023) presented a comprehensive analysis of various algorithms, such as time series and deep learning techniques, for stock market predictions and analysis. They discussed the use of methods like linear regression, ARIMA, and other approaches under the computational framework with a robust result. In the context of technical analysis or computational approach, there exist different indicators, like chart patterns, phases, and candle sticks and are accepted as fundamentally significant and efficient. Li and Bastos (2020), Fauzi and Wahyudi (2016)stated that image processing and regression techniques will further enhance these analyses. Their findings indicate that the technical analysis can yield more accurate results for market forecasting and price prediction. Taking into account of effectiveness of the computational approach, this study highlights the importance of applying the computational approach to examine the financial data. This study will apply the integral Fourier Transforms, which are especially employed in the fields of physics and engineering. The Fourier transform is used to break down the time series data, generally on signals, to convert them from the time domain to the frequency domain. This mathematical technique is applied to analyze and represent the signals by breaking them into frequency elements.
@@
Fourier Transform (FT) is a mathematical method for analyzing time series data and decomposing it into frequency components, especially used to analyze the cyclical movements. FT permits spectral analysis, providing a band of time series data and helping to disclose the dominant frequencies and patterns, like seasonal trends or business cycles. The properties of the Fourier Transform allow stacking the signals into the frequency domain, which simplifies the analysis to understand their interactions to form the signal. This mathematical approach has multiple applications comprising signal processing, engineering, and economics, which enhances our understanding of the cyclical patterns and phases Cochran et al. (1967). The mathematical approach of the Fourier Transform and considering its effectiveness in analyzing the cyclical patterns of the time series data will be employed to explore the National Stock Exchange Data.
@@
Here in this project,the we use FFTW library Frigo and Johnson (2005)(@https://www.fftw.org/) within a computational architecture named TARA(@https://rupakmukherjee.github.io/TARA/) Mukherjee and Ganesh (2019)Mukherjee (2019)Mukherjee et al. (2018a)Mukherjee and Ganesh (2018). The TARA simulation framework is a multi-dimensional pseudo-spectral solver to simulate incompressible and weakly compressible fluid or plasma turbulence Mukherjee et al. (2019a). TARA has the flexibility to add higher-order fluid-moment equations with minimal computational overhead Mukherjee et al. (2019c) Mukherjee et al. (2018b). This architecture runs efficiently on massively parallel CPU as well as multi-GPU architectures (Mahapatra et al. (2018) Mukherjee et al. (2019d)). In addition, the performance scales efficiently under MPI as well as OpenMP (shared- and distributed-memory) clusters Biswas et al. (2021) Biswas et al. (2020) Gupta et al. (2019). Here in this work, we extend the TARA diagnostic module and perform an extended form of extended integral transform analysis (Kim & Powers, 1979) of time-series data. We benchmark our new module with existing analytical results and then perform an analysis of NSE data with our new module.
@@ D Proposition
    The behavior of price index of stock market can be decomposed into a periodic nature, with fluctuations on top of it. Thus a pattern analysis technique can lead to the prediction of market price rallies and become crucial for understanding the nature of the market. Peters (1996)Sharma and Wongbangpo (2002) Akar and Başkaya (2011) Cootner (1962) Healy and Palepu (2001)

### v2

@@ A Introduction
The Financial market is extremely reactive to changes in the market environment. A small information on underlying factors will lead to changes in market behavior and the pricing patterns. The information which are directly or indirectly related to the market affects the price movement and is vital for price discovery. Reflection of companies’ performance and overall dynamics of the market on price provides a foundation for the stakeholders to assess trends and try to discover the price to make informed decisions. The traces of price discovery go back to the 1900s when Bachelier talked about the application of probability theory to analyze the price and gain from options through his work “Theory of speculation”. Followed by Fama (1965)(1970)[1][2] laid the foundation for the efficient market hypothesis, arguing that prices are random and follow a random walk. He also identified three forms of market efficiency where past prices, public information, and private information play a role in determining the prices. But it is not possible to access all the available information, and no one is capable of accessing all the information, resulting in a random nature in price fluctuations and patterns. The study by Fama et al (1969)[3] revealed that the market prepares for splits, adjusts the prices accordingly, and is capable of fully incorporating the information, thus supporting the Efficient Market Hypothesis. Laffont & Maskin(1990) [4] investigates whether the Efficient Market Hypothesis (EMH) holds under two conditions: first, where prices do not reveal private information, and second, where prices do. They analyze how large traders mask the information using Perfect Bayesian Equilibrium and comparative statistics. They found that the market reveals the information, but at the same time, the traders do not prefer to have the market as efficient. This indicates that the EHM is not applicable in imperfect competition when a large group of traders hides information, preventing the price from reflecting the private information that they hold. Mitchel & Mulherin (1994)[5] also found that the news reported in Dow Jones & Company newspaper is positively correlated to the trading volumes and market returns of US securities, indicating a direct relation to the trading activity. Chan et al.(2008)[6] studied information asymmetry between foreign and domestic shares in the Chinese stock market. Their analyses suggested the presence of information asymmetry between foreign and domestic shares in the Chinese stock market. The study by Urquhart&McGroarty (2016)[7] found the traces of an efficient market on S&P 500 and EUROSTOXX 50 by applying AR-GARCH and Variance. Due to the impact of information on price discovery and fluctuations., Fama’s theory of Efficient Market Hypothesis is still relevant and is important to analyze the stages of market patterns and fluctuations to discover the quantitative aspect of the information in terms of price dynamics.
@@
The prices are impacted by all those direct or indirect factors, signaling that price fluctuations emerge from the changes and information in the market. Fama urged that the market incorporates all information in an efficient manner and the market are always efficient[2]. He also stated that the prices are always random and are capable of reflecting all available information, and the pricing patterns are always unpredictable. Due to the importance of the role of information on price discovery and pricing patterns, it makes investors difficult task to predict the market with certainty. But the development of different theories like Arbitrage Pricing Theory(APT) and Capital Assets Pricing Models (CAPM) helps to understand how various factors influence the asset pricing, providing an advantage to the arbitrage [8][9]. Considering the information and the underlying factors, investors try to assess such trends and movements of price to make informed decisions, as such price reflects the overall performance of the listed companies and overall market dynamics. The rising stock prices indicate enhanced company performance, while declining prices indicate downward trends. Stock prices display cyclical movements that are influenced by the dynamic nature of micro as well as macroeconomic factors. followed by the reactions of investors to any economic factors can lead to variations in share prices[10]. Each stock price exhibits unique cyclical patterns, which are studied separately or with the market. It is also important to observe that the cyclical patterns and behaviors of prices may have emerged from trends and autocorrelation[11]. To understand the market dynamics, one needs to analyze the pricing patterns and their cyclical fluctuations, which ultimately help to predict future price movements. Examining the cyclical components of time series data is important for understanding the market trends and forecasting the prices, supporting the interpretation of the movements, whether short or long period[12]. Forecasting the market depends on identifying the patterns in price sequences by deeply analyzing the historical prices to improve predictability[13].
@@
while discussing the price discovery. It is also equally important to recognize and understand that the patterns are not always uniform and are always influenced by the uncertainties and asymmetric information. Such interaction of the elements has offered the interest of examining the price assessment and prediction, ultimately leading an informed decision-making. A study by Malkiel (1989) [14] suggested that to have more research on the dynamics of pricing to discover the true nature of pricing patterns, whether they are truly because of the nature of the market or a result of data exploitation. The fluctuations in the stock market are not always consistent and are impacted by various related factors, as well as events and information asymmetry, which causes turmoil in the market, signaling an unpredictable nature of the price movements. These regular ups and downs of the stock prices and continuous fluctuations of the movement of price are known as turbulence [15] [16]. Such situations in the market create volatility, making it difficult to understand the cyclical patterns and movements of prices. The turbulence in the financial market is generally caused by the situation of information asymmetry and high-energy events, and it is important to analyze the market turbulence and related factors. Engel et al.(2018) [17] studied the major financial crisis, which caused the turbulence in the stock market and framed an early Warning signal using Markov switching models on such situations for Asian and European stock markets. Another study by Khoojine & Han (2020)[18] developed the SPNAR stock prices network auto-regressive model to analyze the Chinese stock market during pre- and post-turbulence. Understanding such situations of uncertainty is essential for market analysis and helps to prevent such situations. Due to the flow of asymmetric information and the dynamic nature of the market, the market becomes volatile with rapid fluctuations, which makes it difficult to analyze and predict the market through the price pattern. So the examination of cyclical pattern ( rapid ups and downs) of the market is valuable to understand the nature of the market as turbulent or not. In this paper, we will analyze the Indian stock market, especially the National Stock Exchange of India, considering the factor of efficient market and the price dynamics through which the nature and state of turbulence within the NSE can be understood. The concepts and techniques of computational physics, especially the spectrum analysis, will be employed to understand the above-mentioned scenario, contributing a different approach for analyzing the market dynamics.
@@ A.1 Stock Market and Turbulence
The stock market is impacted by unpredictable events and the turbulence caused by such events needs to be understood to predict and understand the impact of such events on the stock market. The stock market is unpredictable and due to the rapid changes in environment and the events which affects the financial market leads to disturbances and situation of volatility and such unpredictable ups and downs of the stock market is termed to be known as the state of Turbulence. Turbulence in the stock market refers to rapid and significant increase and decrease in the prices, followed by the situation of unpredictability in the financial market. Turbulence signifies the span of extreme volatility and instability in the financial market which are characterized by the fluctuations and huge level of uncertainty [15]. The word turbulent is commonly used as the fluctuations in price and patterns which replicate the velocity fluctuations in the turbulence [16].
@@
Turbulence can be viewed as a combination of various independent cycles/frequencies present in a system. Typically, in fluid turbulence, a fluid medium can sustain several independent frequencies at the same time. This happens as fluids cannot sustain strong shear forces and fluid elements try to thermalise through conduction as well as convection mechanism. When energy is pumped into a fluid medium, typically the energy resides in large wavenumbers or in long-time cycles of the fluid’s motion. However, this energy cascades into shorter wavenumbers or in short-time cycles soon. This cascading of energy into shorter scales is called the inertial range of turbulence. Finally, the energy is drained out in the form of heat, through tiny viscous scales of the system. Thus, typically a fully-developed turbulent fluid contains three clear scales – the first being the scale at which the energy is being injected into the system, the second is the inertial scale where the energy cascades to the shorter scales and the third is the viscous scale where the energy gets drained out from the medium in the form of heat/radiation. The second phase, that is, the inertial scale, plays the key role in developing the turbulence in a fluid. This inertial scale transfers the energy (/amplitude in the stock market) from large-scale vertices (/long-time cycles) to smaller vertices (/short-time cycles or fluctuations) through some nonlinear interactions. It has been found that in a fully developed turbulence, these nonlinear interactions between different modes (/cycles) actually retain information from both the amplitude as well as the phase of each of the modes (/cycles). Thus, merely an analysis using the energy spectrum falls short in understanding the actual nature of the turbulence present in the system. For example, inside a fully developed turbulence, we expect there will be no correlation between the individual phases, and thus each cycle will have its own unique phase information. In addition, if there is a huge correlation between the phases of individual cycles, one can conclude that the medium has not reached a fully developed turbulent phase. In this report, we describe a method to dig out this phase information from stock market data (using an integral transform method) and check for the quality or degree of turbulence in these individual data. This method is already in use to identify and quantify the level of turbulence in fluids as well as in plasmas [19] [20]. However, to the best of our knowledge, this is the first time we employ this technique to categorize turbulence in stock market data. [21][22]
@@ A.2 Market dynamics and Computational Approach
Due to the effectiveness of fundamental analysis to examine the stock market for long-term investors, it lacks effectiveness on short-term predictions as it considers a range of variables, including macroeconomic factors, industry dynamics, and company-specific elements, to determine the value of financial assets. However, this analysis can be complex, requiring consideration of multiple variables and significant time investment, making it more suitable for long-term investors. [23], [24]. In contrast, due to the desire for quicker gains and faster analysis of market trends, technical analysis emerged. This approach posits that all available information is already reflected in stock prices, allowing analysts to focus on price trends using historical data to capitalize on market movements [25], [26], [27], [28], [29]. Technical analysis has experienced significant growth, particularly among short-term traders. Due to the complexity, noisy, chaotic, volatile, and vast amount of data that needs to be analyzed, machine-learning techniques have emerged and delivered superior results. The computational approach, generally a technical analysis, provides more meticulous outcomes and the ability to analyze big data. The data analysts nowadays are using computational and machine learning to analyze financial data. The stock market data are very complicated and non-stationary, making the traditional approaches typically inaccurate, and they treat stock data as a linear time series. The study by Kumar et al.(2021)[30] considered such issues and reviewed the recent developments on the application of computational methods in the financial markets. Their study found that methods like Artificial Neural Networks, Support Vector Machines are intensively used in place of the traditional approach of data analysis. Similarly, Dunis et al. (2014) [31] also provided different computational techniques that are used for analyzing the stock data. Sonkavde et al. (2023)[32] presented a comprehensive analysis of various algorithms, such as time series and deep learning techniques, for stock market predictions and analysis. They discussed the use of methods like linear regression, ARIMA, and other approaches under the computational framework with a robust result. In the context of technical analysis or computational approach, there exist different indicators, like chart patterns, phases, and candlesticks, which are accepted as fundamentally significant and efficient. Li&Bastos(2020)[33] and Fauzi&Wahyudi(2016) [34]stated that image processing and regression techniques will further enhance these analyses. Their findings indicate that the technical analysis can yield more accurate results for market forecasting and price prediction. Taking into account of effectiveness of the computational approach, this study highlights the importance of applying the computational approach to examine the financial data. This study will apply the integral Fourier Transforms, which are especially employed in the fields of physics and engineering. The Fourier transform is used to break down the time series data, generally on signals, to convert them from the time domain to the frequency domain. This mathematical technique is applied to analyze and represent the signals by breaking them into frequency elements.
@@
Fourier Transform (FT) is a mathematical method for analyzing time series data and decomposing it into frequency components, especially used to analyze the cyclical movements. FT permits spectral analysis, providing a band of time series data and helping to disclose the dominant frequencies and patterns, like seasonal trends or business cycles. The properties of the Fourier Transform allow stacking the signals into the frequency domain, which simplifies the analysis to understand their interactions to form the signal. This mathematical approach has multiple applications comprising signal processing, engineering, and economics, which enhances our understanding of the cyclical patterns and phases [35]. The mathematical approach of the Fourier Transform and considering its effectiveness in analyzing the cyclical patterns of the time series data will be employed to explore the National Stock Exchange Data.
@@
Here in this project,the we use FFTW library [36](@https://www.fftw.org/) within a computational architecture named TARA(@https://rupakmukherjee.github.io/TARA/) [37][38][39][40]. The TARA simulation framework is a multi-dimensional pseudo-spectral solver to simulate incompressible and weakly compressible fluid or plasma turbulence [41]. TARA has the flexibility to add higher-order fluid-moment equations with minimal computational overhead [42] [43]. This architecture runs efficiently on massively parallel CPU as well as multi-GPU architectures [44] [45]. In addition, the performance scales efficiently under MPI as well as OpenMP (shared- and distributed-memory) clusters [46] [47] [48]. Here in this work, we extend the TARA diagnostic module and perform an extended form of extended integral transform analysis (Kim & Powers, 1979) of time-series data. We benchmark our new module with existing analytical results and then perform an analysis of NSE data with our new module.
@@ D Proposition
    The behavior of price index of stock market can be decomposed into a periodic nature, with fluctuations on top of it. Thus a pattern analysis technique can lead to the prediction of market price rallies and become crucial for understanding the nature of the market. [49][50] [51] [52] [53]